OpenStudy (ray1998):

Can someone please help me? Fan and medal will be given! Write an explicit formula for the sequence. Find the seventh term. 2, 4, 8, 16,..... Thank you in advance!

OpenStudy (anonymous):

Maybe there is a common ratio, divide adjacent terms, e.g. 4/2 or 8/4

OpenStudy (jaelenecastro1998):

so the repetition method their using is multiply by 2 2*2=4 4*2=8 8*2=16 16*2=32 32*2=64 64*2=128(Answer)

OpenStudy (ray1998):

Okay, thank you guys!

OpenStudy (jaelenecastro1998):

No problem

OpenStudy (kittykatiehunt):

Jalen you only did 6 you need one more

OpenStudy (jaelenecastro1998):

no cause 2 is number one

OpenStudy (ray1998):

So, from now on, when I come to these equations, I need to do that with whatever the first number of the sequence is, with whatever term is given to me?

OpenStudy (kittykatiehunt):

Yes Ik I got that but u only multiplied 6 of them you need one more multiply the 64 by 2 and that would be the answer

OpenStudy (jaelenecastro1998):

yes first figure what pattern that they are using

OpenStudy (ray1998):

No @Kittykatiehunt , they got it right :)

OpenStudy (ray1998):

Okay! And thanks again :)

OpenStudy (kittykatiehunt):

Alright

OpenStudy (jaelenecastro1998):

No problem